OpenStudy (lncognlto):

How do I work out the least value of 15 cos x - 9 sin x as x varies?

OpenStudy (anonymous):

write it as a single function of sine or cosine

OpenStudy (anonymous):

\[a\cos(x)+b\sin(x)=r\sin(x+\theta)\]

OpenStudy (anonymous):

\(r=\sqrt{a^2+b^2}\) and although you can find \(\theta\) if you want, it is unnecessary for this problem you only need \(r\)

OpenStudy (lncognlto):

Okay, so then it is \[3\sqrt{34}\cos(x+a)\]

OpenStudy (lncognlto):

And so the smallest value of cos would be -1, so the least value would be \[-3\]

OpenStudy (lncognlto):

*-3 (sqr 34)

OpenStudy (anonymous):

yes

OpenStudy (lncognlto):

\[-3\sqrt{34}\]

OpenStudy (lncognlto):

There

OpenStudy (lncognlto):

Cool, thanks, I got it. :)

OpenStudy (anonymous):

yw